OpenStudy (anonymous):

What is he domain of the function √x(x-5) ?

OpenStudy (anonymous):

x>0 \[x \ge 0\]

sam (.sam.):

\[x \in \mathbb{R} : x \geq 0\]

sam (.sam.):

all non-negative real numbers

OpenStudy (anonymous):

may u want it:  x<=0,x>=5